A moving particle encounters an external electric field that decreases its kinetic energy from 9330 eV to 7410 eV as the particle moves from position A to position B. The electric potential at A is -45.0 V, and that at B is +17.0 V.

Determine the charge of the particle.

Include the algebraic sign (+ or -) with your answer.
1
Expert's answer
2020-09-14T12:54:29-0400

from conservation of energy


"\Delta k=\Delta k_B-\Delta k_A=-q\Delta V=-q(V_b-V_A)"


where

K= kinetic energy(at and and at B)

q= charge

V= potential(at A and at B)


therefore making q subject of the formula


"q=-\frac{K_B-K_A}{V_b-V_a}=\frac{7410eV-9330eV}{17-(-45)}=-\frac{-1920eV}{62}=30.968eV"


"30.968eV=30.968e\times\frac{1.6\times 10^{-19}c}{e}"


"q=4.9549\times10^{-18}c"